# Multilingual large language models leak human stereotypes across language boundaries

### v1

@@ Abstract.
Multilingual large language models have been increasingly popular for their proficiency in comprehending and generating text across various languages. Previous research has shown that the presence of stereotypes and biases in monolingual large language models can be attributed to the nature of their training data, which is collected from humans and reflects societal biases. Multilingual language models undergo the same training procedure as monolingual ones, albeit with training data sourced from various languages. This raises the question: do stereotypes present in one social context leak across languages within the model? In our work, we first define the term “stereotype leakage” and propose a framework for its measurement. With this framework, we investigate how stereotypical associations leak across four languages: English, Russian, Chinese, and Hindi. To quantify the stereotype leakage, we employ an approach from social psychology, measuring stereotypes via group-trait associations. We evaluate human stereotypes and stereotypical associations manifested in multilingual large language models such as mBERT, mT5, and ChatGPT. Our findings show a noticeable leakage of positive, negative, and non-polar associations across all languages. Notably, Hindi within multilingual models appears to be the most susceptible to influence from other languages, while Chinese is the least. Additionally, ChatGPT exhibits a better alignment with human scores than other models.
@@ 1. Introduction
Recently, the deployment of large language models has the potential to exacerbate the issue. Large language models are becoming increasingly language-agnostic. For instance, models like ChatGPT(OpenAI111https://openai.com/chatgpt, we use GPT text-davinci-003 version(Ouyang et al., 2022)) and mBART (Lin et al., 2022) can operate without being restricted to a specific language, handling input and output in multiple languages simultaneously. This thus gives rising opportunities for what we refer to as stereotype leakage, or the transmission of stereotypes from one culture to another.
Stereotype leakage within large language models may export harmful stereotypes across cultures and reinforce Anglocentricism222Anglocentrism is the practice of viewing and interpreting the world from an English-speaking perspective with the prioritization of English culture, language, and values. Anglocentrism can lead to biases and neglect of global perspectives and experiences.. Previous works (e.g., Goldstein et al., 2023; Weidinger et al., 2021) have highlighted the potential for language model outputs to change users’ perceptions and behaviors. Stereotype leakage from large language models may further entrench existing stereotypes among model users, as well as create new stereotypes that have transferred from a different language. Therefore, in this work, we investigate the degree of stereotype leakage within multilingual large language models (MLLMs) as a step toward understanding and mitigating stereotype leakage for AI systems.
@@
Large language models are trained on existing language data, and even monolingual language models have been demonstrated to replicate stereotypical associations present in the training data. (Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022). Thus, with the shared knowledge between languages in MLLMs, it is likely that stereotypes may also leak between languages. Stereotypes are abstract and over-generalized pictures drawn about people based on their group membership, and these perceptions can be specific to each culture. Though LLMs are trained on language-based data rather than culture-based data, languages reflect the stereotypes associated with the cultures they represent. Thus, for the purpose of studying stereotypes in MLLMs, we divide the world according to languages, with the understanding that a single language may reflect multiple cultures. Previously, many works have examined Western stereotypes in English language models (e.g. Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022), whereas limited works have attempted to assess stereotypes in multilingual language models (e.g. Kaneko et al., 2022; Levy et al., 2023; Câmara et al., 2022) due to the complexity of stereotypes manifested in various cultures, limited resources, and Anglocentric norms (Talat et al., 2022).
In this paper, we aim to investigate the existence of stereotype leakage in MLLMs, which we define as the effect of stereotypical word associations in MLLMs of one language impacted by stereotypes from other languages. We conduct a human study to collect human stereotypes, adopt word association measurement approaches from previous works (Cao et al., 2022; Kurita et al., 2019) to measure stereotypical associations in MLLMs and analyze the strength and nature of stereotype leakage across different languages both quantitatively and qualitatively.
To test our hypothesis that there are significant stereotypes leaked across languages in MLLMs, we sample four languages: English, Russian, Chinese, and Hindi. We pick languages that come from the Indo-European and Sino-Tibetan language families, ranging from high (English) to low-resource (Hindi) languages333 High-resource languages are languages that have more training data available, while low-resource languages have less.. We measure the degree of stereotype leakage between the four languages in three MLLMs — mBERT, mT5, and ChatGPT444Both the code and the dataset, along with a datasheet (Gebru et al., 2018), are available under a MIT licence at: https://github.com/AnnaSou/Stereotype_Leakage.. Both mBERT and mT5 are back-end MLLMs. MT5 has better multilingual performance than mBERT, whereas mBERT has more comparable monolingual BERT models for the four languages. ChatGPT is one of the state-of-the-art MLLMs that has been popularly deployed to users. With these, we examine the impact of human stereotypes from different languages on stereotypical associations in MLLMs.
@@ 2. Measuring Stereotype Leakage in MLLMs
For each language, we aim to assess the degree of stereotype leakage from the other languages to this target language in MLLMs. Specifically, we measure the effect of human stereotypes from all four languages () on the target language’s MLLM stereotypical association (), as shown in Equation 1. For mBERT, we also measure the impact of the stereotypical association from the target monolingual model (). We use a mixed-effect model to fit the formula and calculate the effect. If the coefficient of a variable is positive and has a p-value of less than , then the variable has a significant effect on . If there are significant effects from the non-target language’s human stereotypes, then there are potential stereotype leakages from this language to the target language.
@@ 2.1. Stereotype Measurement
For the groups, we picked groups listed in Table 2: shared groups with shared stereotypes (groups that are present in all four countries and are expected to be targeted by similar stereotypes), shared groups with non-shared stereotypes (groups that are present in all four countries but expected to be targeted by dissimilar stereotypes), and not shared groups (groups that exist uniquely in each country; three groups for each country). For shared groups, we manually selected groups from the list of social groups from (Cao et al., 2022). To collect not shared groups, we conducted a survey among native speakers. For each language, we asked native speakers to list social groups that they believe are unique to their culture. We then chose social groups per language based on the outcome of the majority vote.
In our human study, we further verified that each group matches the property of its category. To illustrate, stereotypes of groups in the first category exhibit an average correlation score of across languages. In contrast, groups in the second and third categories demonstrate progressively lower correlation scores of and , respectively.
| Agency | powerless | powerful | Beliefs | Communion | untrustworthy | trustworthy | ||
| low status | high status | religious | science-oriented | dishonest | sincere | |||
| dominated | dominating | conventional | alternative | cold | warm | |||
| poor | wealthy | conservative | liberal | benevolent | threatening | |||
| unconfident | confident | traditional | modern | repellent | likable | |||
| unassertive | competitive | egotistic | altruistic | |||||
@@
Shared/ Shared |
man, woman, gay, lesbian, single mother, housewife, software engineer, wealthy person, poor person, disabled person |
Shared/ Non-shared |
Asian person, Black person, Muslim person, immigrant, government official, civil servant, feminist, veteran |
Non-shared/ Non-shared |
USA: Texan, Mormon, Puerto Rican |
@@ 2.1.1. Human stereotypes
To collect human stereotypes, we conducted a human study on Prolific555https://www.prolific.co/ for each of the four languages with native speakers of the respective languages who lived or still live in the United States, Russia, China, and India666Approved by our institutional IRB, #1724519-3.. In the survey, participants were first asked to mark at least social groups that they feel they are familiar with. Then they were asked to rate the group-trait associations of social groups from their list of familiar groups. All surveys were in the respective languages translated by native speakers. For shared/shared and shared/non-shared groups, we collected at least participants’ annotations per group per language. For non-shared groups with non-shared stereotypes, we collected at least annotations for the language they originate from, with no minimum limit of annotations for other languages. See Appendix Appendix A for the survey details and participant demographics.
@@ 2.1.2. Model stereotypical associations
To measure stereotypical group-trait associations in large language models, we adopted different approaches for different MLLMs. For mT5, we used the increased log probability score (ILPS) (Kurita et al., 2019), which computes the likelihood for the model to generate a trait given a templated sentence about a group. For example, indicates the correlation between competent and “Asian people”.
For mBERT, we used the sensitivity test (SeT) (Cao et al., 2022), which is shown to have better alignment with human stereotypes (Cao et al., 2022). It measures how much the model weights would have to change in order to have the trait be the most likely generated words given a templated sentence about a group. SeT captures the model’s confidence in predicting the trait given the group.
For ChatGPT, since the model features are not public, we did not have access to compute the ILPS or SeT scores. Thus, we proposed a new method of measuring stereotypical group-trait associations in GPT-sorted models. We queried ChatGPT for each trait-group pair with the prompt, “You need to write a story about a “woman”. Please, choose either powerless or powerful as the theme of the story.” We repeated the process times for each trait-group pair and counted how many times out of the model picked one trait. This allowed us to measure ChatGPT confidence in the studied traits.
@@ 3. Stereotype Leakage and Its Effects
In this section, we introduce our quantitative and qualitative results of the assessment of stereotype leakage across languages in MLLMs. We study the extent to which human stereotypes from the four languages are represented in the respective languages in MLLMs’ stereotypical associations.
@@ 3.1. Quantitative Results
Based on the measure from Equation 1, we can compute the amount of stereotype leakage from any language to any other language across three models. We present this as a flow in Figure 2, which visualizes for each human source language and model target language how much of the stereotypical associations in the target language model are influenced by human stereotypes in the culture associated with the source language. For instance, in Figure 2, we can see that in ChatGPT stereotypical associations in the Russian language (target language) are influenced by stereotypes observed in human surveys in two source languages: Chinese and Russian. While having the influence from the same language is expected, the leakage happening from the Chinese language is undesirable.
In our analysis of mBERT, we evaluated the influence of monolingual and found that it exerts a stronger impact compared to human stereotypes. Table 3 illustrates that, within mBERT, the Hindi language exhibits the least susceptibility to the influence of monolingual Hindi , while monolingual English BERT demonstrates the strongest influence on the English language within mBERT. This is expected, as English models are the most frequently used and more advanced than models for other languages. Regarding human stereotypes, we observed a significant leakage of stereotypes from Hindi to English and Chinese with coefficients of () and (), respectively. We observed that English human stereotypes manifest in mBERT Hindi with a coefficient of (). Secondly, within the mT5 model, we observed significant stereotypes leakages. Some contributions from Russian and Chinese languages to Hindi were observed. For ChatGPT, we observed significant stereotypes leakages across languages. In terms of intensity, the Russian language has the largest flow from Chinese, which corresponds to a coefficient of (). We also observe a significant impact from English to ChatGPT Hindi with a coefficient of (). We may also see that ChatGPT is the model that is the most affected by human stereotypes encompassing both stereotype leakages and stereotypes originating from the target language itself. Notably, the most significant effects are from human stereotypes of the target languages, which is expected.
Overall, as presented in Figure 2, Hindi is the language that endures the most stereotype leakage – it has cases of significant stereotype leakage from other languages across models. Since Hindi is the only low-resource language we tested, this might explain why it absorbs stereotypes from other languages. The Chinese language has leakages across the models, but in both cases, this comes from the Hindi language. English and Russian languages each have just significant leakage. The first might be explained by the fact that all models were initially trained in English.
| mBERT | mT5 | ChatGPT | |
@@ 3.2. Qualitative Results
Next, we examine the specific stereotypical associations that leak from one language to another and consider the potential influence of such strengthened associations. We focused on the ChatGPT model because it is more influenced by human stereotypes. For each source-target language pair and each group, we looked into the group’s most associated traits from ChatGPT of the target language which were not rated as associated with the group according to human stereotypes of the target language, but match with human stereotypes of the source language. We observed two main types of leakages, wherein positive and negative representations become stronger for certain languages. In other words, we see the leakage of negative stereotypes while there are also cases when some groups acquire more positive representation. In addition, we observed a non-polar leakage, which refers to neither positive nor negative representations.
@@ 3.2.1. Positive Leakage
According to human annotation, “poor people” are more positively perceived in Russian and Hindi languages than in English. We observe the strengthening of such traits as altruistic, sincere, likable for English. “Housewives” become more warm, sincere, trustworthy in English following leakage from Russian and Hindi.
Another example is “immigrants”. Based on human data, we found that people surveyed in Chinese view this group quite favorably since the majority of immigrants to China are highly qualified professionals (Pieke, 2012). We observed the strengthening of such traits as wealthy, rational, sincere, benevolent in Hindi, Russian, and English. Note that in India, this group is not common, as out of annotators only people choose this group as familiar to them. In addition, we observed the strengthening of powerful, trustworthy, and sincere traits for “Asian people” group in Russian and Hindi leaked from Chinese. Another example of the leakage of positive perceptions is for “gay men” and “lesbians” from English to other languages. Such traits as powerful, likable, confident, sincere, trustworthy become stronger. In addition, “Black people” become more strongly associated with the trait liberal in Hindi.

### v2

@@ Abstract.
Multilingual large language models have been increasingly popular for their proficiency in processing and generating text across various languages. Previous research has shown that the presence of stereotypes and biases in monolingual large language models can be attributed to the nature of their training data, which is collected from humans and reflects societal biases. Multilingual language models undergo the same training procedure as monolingual ones, albeit with training data sourced from various languages. This raises the question: do stereotypes present in one social context leak across languages within the model? In our work, we first define the term “stereotype leakage” and propose a framework for its measurement. With this framework, we investigate how stereotypical associations leak across four languages: English, Russian, Chinese, and Hindi. To quantify the stereotype leakage, we employ an approach from social psychology, measuring stereotypes via group-trait associations. We evaluate human stereotypes and stereotypical associations manifested in multilingual large language models such as mBERT, mT5, and GPT-3.5. Our findings show a noticeable leakage of positive, negative, and non-polar associations across all languages. Notably, Hindi within multilingual models appears to be the most susceptible to influence from other languages, while Chinese is the least. Additionally, GPT-3.5 exhibits a better alignment with human scores than other models.
@@ 1. Introduction
Recently, the deployment of large language models has the potential to exacerbate the issue. Large language models are becoming increasingly language-agnostic. For instance, models like ChatGPT111https://openai.com/chatgpt and mBART (Lin et al., 2022) can operate without being restricted to a specific language, handling input and output in multiple languages simultaneously. This thus gives rising opportunities for what we refer to as stereotype leakage, or the transmission of stereotypes from one culture to another.
Stereotypes are abstract and over-generalized pictures drawn about people based on their group membership, and these perceptions can be specific to each culture. Stereotype leakage within large language models may export harmful stereotypes across cultures and reinforce Anglocentricism222Anglocentrism is the practice of viewing and interpreting the world from an English-speaking perspective with the prioritization of English culture, language, and values. Anglocentrism can lead to biases and neglect of global perspectives and experiences.. Previous works (e.g., Goldstein et al., 2023; Weidinger et al., 2021) have highlighted the potential for language model outputs to change users’ perceptions and behaviors. Stereotype leakage from large language models may further entrench existing stereotypes among model users, as well as create new stereotypes that have transferred from a different language. Therefore, in this work, we investigate the degree of stereotype leakage within multilingual large language models (MLLMs) as a step toward understanding and mitigating stereotype leakage for AI systems.
@@
Large language models are trained on existing language data, and even monolingual language models have been demonstrated to replicate stereotypical associations present in the training data. (Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022). Thus, with the shared knowledge between languages in MLLMs, it is likely that stereotypes may also leak between languages. Though LLMs are trained on language-based data rather than culture-based data, languages reflect the stereotypes associated with the cultures they represent. For the purpose of studying stereotypes in MLLMs, we divide the world according to languages, with the understanding that a single language may reflect multiple cultures. Previously, many works have examined Western stereotypes in English language models (e.g. Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022), whereas limited works have attempted to assess stereotypes in multilingual language models (e.g. Kaneko et al., 2022; Levy et al., 2023; Câmara et al., 2022) due to the complexity of stereotypes manifested in various cultures, limited resources, and Anglocentric norms (Talat et al., 2022).
In this paper, we investigate the existence of stereotype leakage in MLLMs. We define stereotype leakage as the effect of stereotypical word associations in MLLMs of one language impacted by stereotypes from other languages. We conduct a human study to collect human stereotypes, adopt word association measurement approaches from previous works (Cao et al., 2022; Kurita et al., 2019) to measure stereotypical associations in MLLMs, and analyze the strength and nature of stereotype leakage across different languages both quantitatively and qualitatively.
To test our hypothesis that there is significant stereotype leakage across languages in MLLMs, we sample four languages: English, Russian, Chinese, and Hindi. We pick languages that come from the Indo-European and Sino-Tibetan language families, ranging from high (English) to low-resource (Hindi) languages333 High-resource languages are languages that have more training data available, while low-resource languages have less.. We measure the degree of stereotype leakage between the four languages in three MLLMs: mBERT, mT5, and GPT-3.5444Both the code and the dataset, along with a datasheet (Gebru et al., 2018), are available under a MIT licence at: https://github.com/AnnaSou/Stereotype_Leakage.. Both mBERT and mT5 are back-end MLLMs. MT5 has better multilingual performance than mBERT, whereas mBERT has more comparable monolingual BERT models for the four languages. GPT-3.5 is one of the state-of-the-art MLLMs that has been popularly deployed to users. With these, we examine the impact of human stereotypes from different languages on stereotypical associations in MLLMs.
@@ 2. Measuring Stereotype Leakage in MLLMs
For each language, we aim to assess the degree of stereotype leakage from the other languages to this target language in MLLMs. Specifically, we measure the effect of human stereotypes from all four languages () on the target language’s MLLM stereotypical association (), as shown in Equation 1. We also control the impact of the stereotypical association from the target monolingual model (). However, since we can only find monolingual BERT model for all four languages, we use these as proxies of for all MLLMs. We use a mixed-effect model to fit the formula and calculate the effect. If the coefficient of a variable is positive and has a p-value of less than , then the variable has a significant effect on . If there are significant effects from a non-target language’s human stereotypes, then there is potential stereotype leakage from this non-target language to the target language.
@@ 2.1. Stereotype Measurement
For the groups, we picked groups listed in Table 2: shared groups with shared stereotypes (groups that are present in all four countries and are expected to be targeted by similar stereotypes), shared groups with non-shared stereotypes (groups that are present in all four countries but expected to be targeted by dissimilar stereotypes), and non-shared groups (groups that exist uniquely in each country). For shared groups, we manually selected groups from the list of social groups from (Cao et al., 2022). To collect non-shared groups, we conducted a survey among native speakers. For each language, we asked native speakers to list social groups that they believe are unique to their culture. We then chose social groups per language based on the outcome of the majority vote.
In our human study, we further verify that each group matches the property of its category. To illustrate, stereotypes of groups in the first category exhibit an average correlation score of across languages. In contrast, groups in the second and third categories demonstrate progressively lower correlation scores of and , respectively.
| Agency | powerless | powerful | Beliefs | Communion | untrustworthy | trustworthy | ||||
| low status | high status | religious | science-oriented | dishonest | sincere | |||||
| dominated | dominating | conventional | alternative | cold | warm | |||||
| poor | wealthy | conservative | liberal | benevolent | threatening | |||||
| unconfident | confident | traditional | modern | repellent | likable | |||||
| unassertive | competitive | egotistic | altruistic | |||||||
@@
| Shared/ Shared | man, woman, gay, lesbian, single mother, housewife, software engineer, wealthy person, poor person, disabled person |
| Shared/ Non-shared | Asian person, Black person, Muslim person, immigrant, government official, civil servant, feminist, veteran |
| Non-shared/ Non-shared | USA: Texan, Mormon, Puerto Rican |
@@ 2.1.1. Human stereotypes
To collect human stereotypes, we conduct a human study on Prolific555https://www.prolific.co/ for each of the four languages with native speakers of the respective languages who lived or still live in the United States, Russia, China, and India666Approved by our institutional IRB, #1724519-3.. In the survey, participants are first asked to mark at least social groups that they feel they are familiar with. Then they are asked to rate the group-trait associations of social groups from their list of familiar groups. All surveys are in the respective languages translated by native speakers. For shared/shared and shared/non-shared groups, we collect at least participants’ annotations per group per language. For non-shared groups with non-shared stereotypes, we collect at least annotations for the language they originate from, with no minimum limit of annotations for other languages. See Appendix A for the survey details and participant demographics.
@@ 2.1.2. Model stereotypical associations
To measure stereotypical group-trait associations in large language models, we adopt different approaches for different MLLMs. For mT5, we use the increased log probability score (ILPS) (Kurita et al., 2019), which computes the likelihood for the model to generate a trait given a templated sentence about a group. For example, indicates the correlation between competent and “Asian people”.
For mBERT, we use the sensitivity test (SeT) (Cao et al., 2022), which is shown to have better alignment with human stereotypes (Cao et al., 2022). It measures how much the model weights would have to change in order to have the trait be the most likely generated words given a templated sentence about a group. SeT captures the model’s confidence in predicting the trait given the group.
In evaluating GPT-3.5, we analyze the model in a generative context, which is its designated use case. For each group, we measure its association with various traits through pairwise comparisons. We prompt the model to generate a story about the group centered around one of the traits from the trait pairs. We utilize the following prompt in English and request translations into Russian, Chinese, and Hindi from native speakers to assess the model across different languages. For each group and trait pair, we query the model ten times and compute the normalized probability score for each trait. Subsequently, we average the trait scores from all trait pair comparisons to establish the group-trait association score. Our experimentation employs the gpt-3.5-turbo-0125 model. Notably, we observe a tendency in the GPT model to favor positive traits (e.g., powerful, trustworthy) over negative traits (e.g., powerless, untrustworthy). To ensure equitable comparisons, we juxtapose positive traits with positive ones and negative traits with negative ones.
prompt sys = “You are a writer.”
prompt = “You are writing about [GROUP]. Before writing, think about what theme you want to pick. You can choose either ‘[TRAIT1]’ or ‘[TRAIT2]’ as your theme. You can also choose ‘neither’ if you think neither of these themes fits. Note that you can choose only one theme. Output the exact name of the theme only, without any punctuation.”
When processing the outputs of GPT-3.5, we employ an exact match criterion to assign scores to traits. For traits comprising sub-tokens, we sum the log probabilities of the sub-tokens to determine the trait’s score. In instances where outputs do not precisely match the traits in the prompt, we have them manually processed by native speakers of the respective language. Throughout this process, we consistently observe system failures or the generation of stereotypical outputs for marginalized groups. Notably, for certain groups like “feminist” and “Muslim person” in Chinese, the model often disregards the prompt and simply outputs the group name. Moreover, in some cases, the model alters the trait specified in the prompt. For example, it changes dominating to dominated for “disabled person” in English or poor to wealthy for “migrant worker” in Russian. Additionally, the model may overlook the traits provided in the prompt and generate stereotypical traits instead. For instance, in Russian, it generates rape and patriot for “Puerto Rican” or cowboy for “Texan”. These occurrences can potentially cause both representational and quality of service harm to stakeholders of the model. While we do not explicitly analyze these patterns, we believe it is imperative for future research to thoroughly investigate them.
@@ 3. Stereotype Leakage and Its Effects
In this section, we present our quantitative and qualitative results of the assessment of stereotype leakage across languages in MLLMs. We study the extent to which human stereotypes from the four languages are represented in the respective languages in MLLMs’ stereotypical associations.
@@ 3.1. Quantitative Results
We compute the stereotype leakage across languages within three MLLMs based on Equation 1. The findings are presented in Figure 2, illustrating the extent to which stereotypical associations in the target language model are influenced by human stereotypes present in the culture associated with the source language. For example, in Figure 2, we observe that within GPT-3.5, stereotypical associations in the English language (target language) are influenced by human stereotypes from two distinct source languages: Russian and Hindi. This observation suggests the presence of stereotype leakage within the GPT-3.5 model.
In our analysis of mBERT, we observe significant leakages of stereotypes from Hindi to English and Chinese with coefficients of () and (), respectively. We also observe English human stereotypes manifesting in mBERT Hindi with a coefficient of (). Within the mT5 model, we find two significant stereotype leakages, both of which are leakages targeting Hindi. Russian and Chinese human stereotypes manifest in mT5 Hindi with coefficients of () and (), respectively. For GPT-3.5, we observe the most significant stereotype leakages across languages, totaling seven. We see most stereotypes leaking from English to all three other languages. The largest flows are from English to Chinese and Hindi, with coefficients of (). Meanwhile, all languages are prone to be affected by leakages from other languages. Overall, we observe that GPT-3.5 is the model most affected by human stereotypes, encompassing both stereotype leakages and stereotypes originating from the target language itself.
Moreover, among all languages, Hindi experiences the highest degree of stereotype leakage — it has four cases of significant stereotype leakage from other languages across three MLLMs. Since Hindi is the only low-resource language we tested, this might explain why it absorbs stereotypes from other languages. Both Chinese and English languages have three leakages across the models. The Chinese language has leakages from all three other languages, while the English language has the most leakage from Hindi. The Russian language has two significant leakages from English and Hindi.
Finally, we report the coefficients of effects from monolingual language models () in Table 3. All the effects are statistically significant and are stronger than the effects from human stereotypes. This is not surprising because monolingual language models and multilingual language models share similar training data and model structures.
| mBERT | mT5 | GPT-3.5 | |
@@ 3.2. Qualitative Results
Moving forward, we delve into the specific stereotypical associations that leak from one language to another, considering the potential implications of such strengthened associations. We focus on the GPT-3.5 model, in which we observe the most leakage from human stereotypes. Within each source-target language pair, which has significant stereotype leakage according to our results above, and for each group, we scrutinize the group’s most associated traits from GPT-3.5 in the target language that are not deemed associated with the group according to human stereotypes of the target language but align with human stereotypes of the source language. Our analysis reveals two primary types of leakages: the amplification of positive and negative representations for certain languages. In other words, we observe the leakage of negative stereotypes, alongside instances where certain groups acquire more positive representations. Additionally, we identify non-polar leakages, characterized by neither positive nor negative representations.
@@ 3.2.1. Positive Leakage
According to human annotation, “Asian people” are more positively perceived in English language than in Russian. We observe the strengthening of such traits in GPT-3.5 Russian language as wealthy, likable, and high status. Moreover, “housewives” become more warm in English following leakages from Russian and Hindi. “Black people” are more powerful, modern, confident, and wealthy in the English language following leakage from Hindi. Another example of the leakage of positive perceptions is for “gay men” and “lesbians” from English to other languages. Traits such as likable, confident, warm, dominant, sincere, and powerful become stronger in Russian, Chinese, and Hindi.